+ | R1 ہی ہے۔ ۸۸۸ ة -(1+ R5 | + ابرها -1)What is the output voltage V0 if op-amp is ideal?  

R1= 2 KΩ, R2=8 KΩ, R3=3.2 KΩ, R4=6 KΩ, R5=19 KΩ, R6=4.6 KΩ, RL=9.6 KΩ, V1=1V, I2=0.5 mA and V3=2.5 V.

2) for the same circuit, R1= 2 KΩ R2=8 KΩ, R3=3.6 KΩ, R4=6 KΩ, R5=16.9 KΩ, R6=15 KΩ, RL=10 KΩ, V1=1V, I2=0.5mA and V3=2V. What is the input impedance as seen by source V3?
Answer #1

Ra No B N- ☺ R1 Rs Va RL Vt A R2 R₂ -2 of LIN , T It -111 1 1 R = 2K1, R2 = 8K1, R = 30261, Rw=6K4, R5 =1989 Ro = 4,6K1, R2 =Vi va by virtual short con oncept Ve= V. -163 -1.63 volt apply kell at rode B V gta v RI va -¿z + Ra Un-vo Ro -1.63-1 2R + 11where R3 = 3.6K4, RUS 6 Km Vz=2V Vrt 2 un + 6 3.6k БК Va(1+0.6) = -2 Val 5-1:25 volt -1.2572 I = 3.6K I = +0-208 m Arup - SH